# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v1

@@ 4.1 Forecasting Result
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.403 0.398 0.387 0.400
0.513
0.496
0.419
0.419
0.400 0.406
0.429
0.425
0.435
0.437
0.403
0.407
0.448
0.452
0.481
0.456
0.588
0.517
ETTm2
0.286 0.327 0.281 0.326
0.757
0.610
0.358
0.404
0.291 0.333
0.293
0.342
0.409
0.436
0.350
0.401
0.305
0.349
0.306
0.347
0.327
0.371
ETTh1
0.441 0.429
0.469
0.454
0.529
0.522
0.541
0.507
0.458
0.450
0.542
0.510
0.491
0.479
0.456 0.452 0.440
0.460
0.570
0.537
0.496
0.487
ETTh2
0.380 0.402 0.387 0.407
0.942
0.684
0.611
0.550
0.414 0.427
0.439
0.452
0.602
0.543
0.559
0.515
0.437
0.449
0.526
0.516
0.450
0.459
Electricity
0.215
0.292 0.205 0.290
0.244
0.334
0.251
0.344
0.192 0.295
0.208
0.323
0.229
0.329
0.212
0.300
0.214
0.327
0.193
0.296
0.227
0.338
Exchange
0.362
0.402
0.259 0.281 0.259 0.315 0.271 0.320
0.416
0.443
0.410
0.427
0.385
0.447
0.354
0.414
0.519
0.500
0.461
0.454
0.613
0.539
Traffic
0.623
0.372 0.367
0.404
0.940
0.707
0.370
0.413
0.620
0.336
0.621
0.396
0.622
0.392
0.625
0.383
0.610
0.376
0.624
0.340
0.628
0.379
Weather
0.260 0.283
0.481
0.304
0.550
0.304
0.760
0.473
0.259 0.287
0.271
0.334
0.261
0.312
0.265
0.317
0.309
0.360
0.288
0.314
0.338
0.382
$1^{\text{st}}$ Count
8 20 16 13
1
0
1
0
7 5
0
1
0
0
1
0
5
0
1
1
0
0
$2^{\text{nd}}$ Count 13 7 7 15
1
3
4
3
6 5
2
1
0
0
2
2
0
0
5
4
0
0
$3^{\text{rd}}$ Count
5 5
4
3
3
3
2
2
11 20
4
0
3
0
6 4
2
1
0
1
0
1
4.2 Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 5. It is worth noting that the order of magnitude of the MSE in Figure 5 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. Full hyperparameter sensitivity experiment results are in the Appendix Table 8.
(a) $lr$ $|S|$MSE
(b) Violin diagram of $|S|$ and MSE
Figure 5: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
Table 3: The variance of model error for traversal learning rate and sample space overhaul
Metric Dataset
($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
4.3 Ablation Study
Table 4: The results of the ablation experiment
Metric Models
Dataset
Electricity
ETTh1
ETTh2
ETTm1
ETTm2
Exchange
Traffic
Weather
MSE C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
MAE C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
4.4 Efficiency Comparison
Figure 6: Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 6 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear [18] performance and almost PatchTST [12] accuracy, making it a performance error balanced model.
4.5 Interpretability Analysis
TEFN will allocate mass based on the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, and mass indicates the confidence reflecting support of different possible events. This support allocation is interpretable because BPA allocation utilizes fuzzy membership $\mu(x)=wx+b$ in Figure 7. Essentially, different fuzzy sets are constructed through TEFN for inference.
(a) Channel BPA
(b) Time BPA
Figure 7: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.6 Limitation and Future Prospects
5 Conclusion
This paper treats different channels and different time points of time series as multiple sources of information. By utilizing evidence theory to construct different basic probability assignments, TEFN discards irrelevant time points and channels. Experimentally, TEFN is a parameter-insensitive and stable prediction model with a small parameter count and fast training speed. In future work, we will explore how to optimize basic probability assignments and information fusion in neural networks.
Acknowledgments
The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078).
Appendix A Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [13]. For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The loss function selected during the experiment was MSE, and the optimizer selected Adam [9]. We used Mean Absolute Error (MAE) and Mean Square Error (MSE) in Equation 9 and Equation 10 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (9)
$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (10)
Appendix B Supplementary Experimental Results
B.1 Dataset Feature
In the experimental section, we mainly referred to the experimental dataset and experimental settings of TimesNet. The experimental settings of TimesNet are comprehensive and reasonable [16]. Table 5 shows the quantitative features of the dataset used in the experiment. Dataset size corresponds to the training set, validation set, and test set. Each dataset has four forecasting tasks of different lengths $L\in\{96,192,336,720\}$.
Table 5: The quantitative features of the dataset
Dataset Dimension Dataset Size
Electricity 321 (18317, 2633, 5261)
ETTm1, ETTm2 7 (34465, 11521, 11521)
ETTh1, ETTh2 7 (8545, 2881, 2881)
Traffic 862 (12185, 1757, 3509)
Weather 21 (36792, 5271, 10540)
Exchange 8 (5120, 665, 1422)
B.2 Forecasting Visualization
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 [1] as an example and selected typical models PatchTST [12], Crossformer [20], TimesNet [16], Dlinear [18] and Stationary [11] for visualization in Figure 8. From the visualization results, TEFN fits the details and trends of time series better.
(a) TEFN (Ours)
(b) PatchTST
(c) Crossformer
(d) TimesNet
(e) Dlinear
(f) Stationary
Figure 8: Visualization of Electricity-96 tasks
B.3 Full Comparative Results
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 10 comparative models in terms of forecasting length $l\in\{96,192,336,720\}$ in Table 7. It can be seen that TEFN has positive predictive performance on the experiment dataset, achieving SOTA in multiple forecasting tasks across multiple datasets.
Table 6: Full results of the ablation experiment
Metric Target Model Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
Table 7: Full results of comparative results
Models
TEFN
PatchTST
Crossformer
TiDE
TimesNet
ETSformer
LightTS
Dlinear
FEDformer
Stationary
Autoformer
[12]
[20]
[5]
[16]
[15]
[19]
[18]
[21]
[11]
[17]
Metric
@@ B.3 Full Comparative Results
1
B.4 Full Ablation Study
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 6. In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 11.
$l_{C+T}\leqslant min(l_{C},l_{T})$ (11)
B.5 Full Hyperparameter Sensitivity
The GPU memory in the experimental environment is limited, so we conducted a 2-dimensional traversal of the sample space size $|S|\in\{0,1,2,3,4,5,6\}$ and learning rate $lr\in\{1*10^{-2},5*10^{-2},1*10^{-1}\}$ in Table 8. In the case of sufficient GPU memory, more hyperparameters can be traversed. Based on the complete results of hyperparameter sensitivity, it can be seen that TEFN is an insensitive model to hyperparameters, which means that random hyperparameters can achieve stable and superior prediction performance. We use variance to describe the fluctuation of hyperparameters in Table 3.
Table 8: Full results of hyperparameter sensitivity

### v2

@@ 4.1. Dataset
0
0
0
Table 3. Full results of hyperparameter sensitivity
@@
4.2. Comparative Baseline
To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models. These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities. The selected baseline models are as follows:
  • Transformer-based Models:
    • iTransformer (ICLR 2024) (Liu et al., 2023): A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
    • PatchTST (ICLR 2023) (Nie et al., 2023): This model divides the time series into patches and applies transformer-based operations on each patch independently. It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
    • Crossformer (ICLR 2023) (Zhang and Yan, 2023): This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
    • FEDformer (ICML 2022) (Zhou et al., 2022): A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
    • ETSformer (Woo et al., 2022): A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
    • Stationary (NeurIPS 2022) (Liu et al., 2022): A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
    • Autoformer (NeurIPS 2021) (Wu et al., 2021): A self-supervised transformer model that learns to forecast future values without relying on explicit labels. It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
  • Linear Models:
    • Rlinear (ICLR 2024) (Li et al., 2023): This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data. It is effective at capturing periodic patterns and works well with longer input sequences.
    • Dlinear (AAAI 2023) (Zeng et al., 2023): A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data. It serves as a simple yet effective baseline for comparison.
    • LightTS (Zhang et al., 2022): A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
  • Other Models:
    • TiDE (ICLR 2023) (Das et al., 2023): A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
    • TimeNet (ICLR 2023) (Wu et al., 2023): A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
By comparing TEFN against this diverse set of baseline models, we can comprehensively assess its performance and identify its unique advantages and contributions to the field of time series forecasting.
4.3. Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch (Paszke et al., 2019). For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The open source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam (Kingma and Ba, 2015). We used Mean Absolute Error (MAE) and MSE in Equation 15 and Equation 16 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
(15) $MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
(16) $MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
4.4. Forecasting Result
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table 2. The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting. TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns. Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer (Liu et al., 2023) and PatchTST (Nie et al., 2023), demonstrating its efficiency and scalability.
TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics. Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer (Liu et al., 2023) and $0.025$ smaller than PatchTST (Nie et al., 2023) in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset (Li et al., 2021). Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer. TEFN was $0.008$ smaller than iTransformer (Liu et al., 2023) and $0.007$ smaller than PatchTST (Nie et al., 2023) in predicting a length of $L_{pred}=96$ on the Exchange dataset (Wu et al., 2021). These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
(a) TEFN (Ours)
(b) PatchTST
(c) Crossformer
(d) TimesNet
(e) Dlinear
(f) Stationary
Figure 5. Visualization of Electricity-96 tasks
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 (ecl, [n.d.]) as an example and selected typical models PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), TimesNet (Wu et al., 2023), Dlinear (Zeng et al., 2023) and Stationary (Liu et al., 2022) for visualization in Figure 5. From the visualization results, TEFN fits the details and trends of time series better.
4.5. Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 6. It is worth noting that the order of magnitude of the MSE in Figure 6 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. Full hyperparameter sensitivity experiment results are in the Table 3.
(a) $lr$ $|S|$MSE
(b) Violin diagram of $|S|$ and MSE
Figure 6. Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
Table 4. The variance of model error for traversal learning rate and sample space
Metric Dataset
($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
(a) Channel BPA
(b) Time BPA
Figure 7. Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.6. Ablation Study
Table 5. The results of the ablation experiment
Metric Target Model Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 5. In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 17.
(17) $l_{C+T}\leqslant min(l_{C},l_{T})$
4.7. Efficiency Comparison
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 8 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear (Zeng et al., 2023) performance and almost PatchTST (Nie et al., 2023) accuracy, making it a performance error balanced model.
Figure 8. Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the official code repository (https://github.com/ztxtech/Time-Evidence-Fusion-Network), TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer (Zhang and Yan, 2023) when deployed on portable computers.
4.8. Interpretability Analysis
The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable. TEFN allocates mass to events within the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, where each event represents a potential outcome or scenario. The mass assigned to an event reflects the level of confidence or support for that event, providing insight into the model’s belief about its likelihood of occurrence. This support allocation is particularly interpretable due to the use of fuzzy membership functions $\mu(x)=wx+b$ within the BPA module, as depicted in Figure 7. By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions. For instance, it becomes evident which features are deemed more important than others, and how the model responds to variations within these features. This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
In summary, TEFN’s interpretability stems from its foundational reliance on fuzzy logic and the Dempster-Shafer theory, coupled with the transparency of its BPA and fuzzy reasoning processes. By exploring these aspects, users can develop a comprehensive understanding of the model’s decision-making process and prediction outcomes, fostering trust and confidence in the model’s capabilities.
4.9. Limitation and Future Prospects
5. Conclusion
This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets. TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data. This selective information fusion significantly reduces the computational burden, enabling TEFN to efficiently process and predict within acceptable time frames. The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.
Acknowledgements.
The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078)